OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

i need help

OpenStudy (skullpatrol):

I can only help you help yourself.

OpenStudy (anonymous):

thats ok thank you can you help me understand the problem

OpenStudy (skullpatrol):

Yes, askaway.

OpenStudy (anonymous):

i just need you to possibly walk me through it so i can do the other three

OpenStudy (skullpatrol):

What do you know about your equation: $$\Huge y = 2x + 2 $$

OpenStudy (anonymous):

that it is really confusing and i need to isolate the variable correct

OpenStudy (skullpatrol):

This equation is in a form called the slope-intercept form, have you heard about that before?

OpenStudy (anonymous):

yes i was not very good at it i am in florida virtual school so no teacher to explain

OpenStudy (skullpatrol):

Have you tried the Khan Academy on YouTube for an explanation?

OpenStudy (anonymous):

i have never heard of it

OpenStudy (skullpatrol):

http://www.youtube.com/watch?v=R948Tsyq4vA

OpenStudy (anonymous):

@AnImEfReaK you ok

OpenStudy (anonymous):

Perpendicular means m1*m2 = -1

OpenStudy (anonymous):

y = 2x + 2 slope = 2 = m1

OpenStudy (anonymous):

2 * m2 = -1 m2 = -1/2 now use two point form u find the equattion

OpenStudy (anonymous):

so its y = −one halfx + 3 or y=2x+6

OpenStudy (anonymous):

i think its one of them

OpenStudy (anonymous):

(6, 3) slope = -1/2 find the equation

OpenStudy (anonymous):

Lines which are perpendicular have their gradients, when multiplied together = -1 In the rule m1 * m2 = -1 m1 is the gradient of the line given = 2 m2 is the gradient of the line we are looking for (which is perpendicular) 2 * m2 = -1 therefore the gradient of the line is m2 = -1/2 Using the general intercept formula y=mx + c y = -1/2x + c We still need to find c Sub in the point (6,3) (y,x) 6= -1/2(3) + c 6= -1.5 + c c = 7.5 Therefore the line is y= -1/2x + 7.5

OpenStudy (anonymous):

thanks give skull patrol a medal plz

OpenStudy (skullpatrol):

Why me?

OpenStudy (anonymous):

u helped me too but i can give only one medal

OpenStudy (skullpatrol):

I can only help you help yourself.